Which has more area a circle 4 inch in diameter or rectangle 4 inch long and 3 inch wide?
Answer by checkley77(12844)

Area of the circle=pir^2=3.14*2^2=3.14^4=12.56 in.^2
Area of the rectangle=4*3=12 in.^2
The circle is .56 in.^2 larger.